# url]marylaisalvana.com

### v1

@@ 1 Introduction
Power grid failures represent an increasingly critical socioeconomic risk as climate change intensifies extreme weather events and infrastructure dependencies deepen. The February 2021 Texas power crisis—which resulted in over 200 deaths and approximately \$130 billion in economic damage—highlighted a critical need for improved predictive capabilities in identifying system-wide vulnerability before catastrophic failures occur [1]. Traditional predictive approaches to grid reliability typically include fragility models [2], spatio-temporal correlation analysis [3, 4], machine learning and deep learning methods [5, 6, 7, 8]. However, these approaches failed to anticipate the catastrophic failure of the Texas power grid in 2021 due to several fundamental limitations. Fragility models generally focus on individual component resilience rather than emergent system-wide behaviors. Correlation-based approaches often assume historical patterns will persist, making them inadequate when faced with unprecedented conditions. Even sophisticated machine learning techniques struggle with the "black swan" nature of such events, as they rely heavily on training data that may not capture the complex interactions and cascading effects that manifest during extreme scenarios.
The inherent difficulty in predicting low-probability, high-consequence events stems from their statistical rarity and the absence of comprehensive historical records. Power systems increasingly operate in regimes for which no precedent exists, driven by changing climate patterns, evolving energy mixes, and shifting demand profiles. The Texas grid failure exemplified this challenge—traditional reliability metrics failed to signal the impending system-wide collapse [9]. Furthermore, traditional methods typically assess risk through deterministic or probabilistic frameworks that struggle to capture how small, seemingly manageable disturbances can trigger disproportionately large cascading failures when systems approach critical states.
This fundamental gap in predictive capability necessitates novel approaches that can detect early warning signals of catastrophic failure in complex systems. One promising direction draws from complexity science, particularly the theory of Self-Organized Criticality (SOC), which offers a framework for understanding how complex systems naturally evolve toward critical states characterized by power-law distributions of event sizes [10, 11]. By analyzing the statistical signatures of SOC in historical outage data, we can potentially identify when power grids approach dangerous thresholds before catastrophic failures materialize [12].
@@ 2.1 The SOC Theory
First proposed by [10], SOC theory describes how complex systems with many interacting components naturally evolve toward critical states—poised between order and disorder—where small perturbations can trigger large-scale cascading events. At criticality, the probability distribution of event sizes follows a power law:
@@
where $x$ represents event size (in our case, number of customers affected by an outage) and $\alpha$ is the critical exponent. This power-law signature has been observed in numerous complex systems, including earthquakes, forest fires, financial markets, and power grids [13, 11].
The critical exponent $\alpha$ serves as a key parameter characterizing system dynamics. When $\alpha>1$, the system maintains a finite expected event size, indicating relative stability despite fluctuations. However, as $\alpha$ approaches and falls below 1, the expected event size diverges, signaling a transition towards a regime where system-wide catastrophic events become increasingly probable [14]. This property makes $\alpha$ a potential leading indicator for critical transitions in complex systems.
Previous work by [12] and [11] identified power-law distributions in historical blackout data, suggesting that power grids may indeed operate as self-organized critical systems. Our approach extends this work by developing a temporal analysis framework to track the evolution of $\alpha$ and evaluate its predictive capability for catastrophic events.
@@ 2.2 Data Collection and Processing
We retrieve outage data from the publicly available Oak Ridge National Laboratory’s Environment for Analysis of Geo-Located Energy Information (EAGLE-I) power outage dataset [15]. EAGLE-I has been collecting U.S. electric outage data since 2014, providing a comprehensive historical record of power disruptions across USA. The dataset includes:
@@ 2.3 Power-Law Parameter Estimation
To estimate the critical exponent $\alpha$, we employed log-linear regression on the frequency distribution of outage sizes using backward-looking temporal windows. For each time point $t$, we analyzed data from the preceding 180-day window (from $t-180$ days to $t$), performing the following steps:
@@ 2.4 Temporal Evolution of System Criticality
The temporal evolution of power system vulnerability can be characterized through systematic tracking of the critical exponent, $\alpha$. Building on theoretical foundations in complexity science [10, 11], we developed a methodology to quantify how system criticality evolves over time:
@@
This approach allowed us to identify critical slowing down and critical fluctuations—hallmark signatures of systems approaching phase transitions [16]. In the context of power grid reliability, a phase transition represents the shift from a regime where outages remain localized and manageable to one where they propagate throughout the system in cascading failures with potentially catastrophic consequences. By tracking the critical exponent’s trajectory relative to the theoretical threshold ($\alpha=1$), we could observe the system’s progression toward criticality, characterized by power-law distributions with increasingly heavier tails.
@@ 2.5 Critical Transition Detection
Complex systems, including power grids, can experience abrupt shifts from one stable state to another, known as critical transitions [16, 17]. The core challenge in complexity science is identifying early warning signals before these transitions occur. Our approach addresses this challenge by analyzing the critical exponent $\alpha$ derived from the power-law distribution of outage sizes.
In SOC theory, the critical exponent $\alpha$ determines the probability distribution of event magnitudes. When $\alpha$ is large, the probability of large-scale events drops rapidly, making them extremely rare. Conversely, when $\alpha$ is small, the decline in probability is more gradual, making larger events relatively more probable. Most importantly, there exists a specific value of $\alpha$ where the system’s behavior fundamentally changes. We identify $\alpha_{c}=1$ as a theoretically significant threshold that separates different regimes of system behavior:
  • Subcritical regime ($\alpha>1$): The system maintains a finite expected value (mean) of event sizes, with limited cascade potential
  • Critical point ($\alpha=1$): The system operates exactly at the boundary between stability and instability
  • Supercritical regime ($\alpha<1$): The system enters a state where extremely large events become increasingly probable and widespread cascading failures can occur
This classification has a precise mathematical foundation based on the expected value of the event size distribution:
[…]
For values of $\alpha$ greater than 1 (subcritical regime), this integral converges to a finite value, indicating that the system maintains a finite mean number of events. However, when $\alpha$ falls below 1 (supercritical regime), the integral diverges, meaning the expected event size becomes mathematically infinite. In practical terms, this signals a regime shift where extremely large events become not just possible but statistically inevitable.
@@
  • The transition from subcritical to supercritical regimes as $\alpha$ crosses the critical point
  • The duration of periods when the system remains in the supercritical regime
  • The depth of excursion into the supercritical regime (how far $\alpha$ drops below 1)
We hypothesize that these three aspects—regime transitions, persistence, and excursion depth—together characterize the vulnerability of power systems to catastrophic failures. Specifically, regime transitions signal the onset of vulnerability, persistence indicates the duration of the vulnerable state, and excursion depth correlates with the potential severity of failures. This theoretical framework provides a foundation for interpreting the critical exponent’s behavior in relation to system-wide risk and offers a quantitative approach to identifying periods of heightened vulnerability before catastrophic failures materialize.
The remainder of this paper applies this framework to analyze the Texas power grid from 2014-2022, with particular attention to the February 2021 crisis. The February 2021 Texas power crisis provides a unique case study for testing these theoretical predictions. By analyzing how the critical exponent evolved through subcritical and supercritical regimes before, during, and after this well-documented catastrophic failure, we can evaluate whether the theoretical framework accurately captures real-world system behavior. This analysis helps establish whether monitoring transitions between critical regimes can provide actionable early warning signals before cascading failures in power systems and potentially other complex infrastructure networks.

### v2

@@ 1 Introduction
Power grid failures represent an increasingly critical socioeconomic risk as climate change intensifies extreme weather events and infrastructure dependencies deepen. The February 2021 Texas power crisis—which resulted in over 200 deaths and approximately \$130 billion in economic damage—highlighted a critical need for improved predictive capabilities in identifying system-wide vulnerability before catastrophic failures occur [txc_winterstorm_2021]. Traditional predictive approaches to grid reliability typically include fragility models [karagiannakis2025fragility], spatio-temporal correlation analysis [lee2023quantifying, bhusal2024estimating], machine learning and deep learning methods [xie2020review, madasthu2023ensemble, yang2023assessing, wang2025deep]. However, these approaches failed to anticipate the catastrophic failure of the Texas power grid in 2021 due to several fundamental limitations. Fragility models generally focus on individual component resilience rather than emergent system-wide behaviors. Correlation-based approaches often assume historical patterns will persist, making them inadequate when faced with unprecedented conditions. Even sophisticated machine learning techniques struggle with the "black swan" nature of such events, as they rely heavily on training data that may not capture the complex interactions and cascading effects that manifest during extreme scenarios.
The inherent difficulty in predicting low-probability, high-consequence events stems from their statistical rarity and the absence of comprehensive historical records. Power systems increasingly operate in regimes for which no precedent exists, driven by changing climate patterns, evolving energy mixes, and shifting demand profiles. The Texas grid failure exemplified this challenge—traditional reliability metrics failed to signal the impending system-wide collapse [busby2021cascading]. Furthermore, traditional methods typically assess risk through deterministic or probabilistic frameworks that struggle to capture how small, seemingly manageable disturbances can trigger disproportionately large cascading failures when systems approach critical states.
This fundamental gap in predictive capability necessitates novel approaches that can detect early warning signals of catastrophic failure in complex systems. One promising direction draws from complexity science, particularly the theory of Self-Organized Criticality (SOC), which offers a framework for understanding how complex systems naturally evolve toward critical states characterized by power-law distributions of event sizes [bak1996nature, dobson2007complex]. By analyzing the statistical signatures of SOC in historical outage data, we can potentially identify when power grids approach dangerous thresholds before catastrophic failures materialize [carreras2004evidence].
@@ 2.1 The SOC Theory
First proposed by [bak1996nature], SOC theory describes how complex systems with many interacting components naturally evolve toward critical states—poised between order and disorder—where small perturbations can trigger large-scale cascading events. At criticality, the probability distribution of event sizes follows a power law:
@@
where $x$ represents event size (in our case, number of customers affected by an outage) and $\alpha$ is the critical exponent. This power-law signature has been observed in numerous complex systems, including earthquakes, forest fires, financial markets, and power grids [sornette2006critical, dobson2007complex].
The critical exponent $\alpha$ serves as a key parameter characterizing system dynamics. When $\alpha>1$, the power-law distribution is normalizable and the system exhibits well-defined probabilistic behavior with statistically rare extreme events. However, as $\alpha$ approaches and falls below 1, the distribution becomes increasingly heavy-tailed, and the relative probability of extreme events grows without bound [pruessner2012self]. This property makes $\alpha$ a potential leading indicator for critical transitions in complex systems.
Previous work by [carreras2004evidence] and [dobson2007complex] identified power-law distributions in historical blackout data, suggesting that power grids may indeed operate as self-organized critical systems. Our approach extends this work by developing a temporal analysis framework to track the evolution of $\alpha$ and evaluate its predictive capability for catastrophic events.
@@ 2.2 Data Collection and Processing
We retrieve outage data from the publicly available Oak Ridge National Laboratory’s Environment for Analysis of Geo-Located Energy Information (EAGLE-I) power outage dataset [brelsford2024dataset]. EAGLE-I has been collecting U.S. electric outage data since 2014, providing a comprehensive historical record of power disruptions across USA. The dataset includes:
@@ 2.3 Power-Law Parameter Estimation
To estimate the critical exponent $\alpha$, we employed log-linear regression on the frequency distribution of outage sizes using backward-looking temporal windows. For each time point $t$, we analyzed data from the preceding 180-day window (from t180t-180 days to $t$), performing the following steps:
@@ 2.4 Temporal Evolution of System Criticality
The temporal evolution of power system vulnerability can be characterized through systematic tracking of the critical exponent, $\alpha$. Building on theoretical foundations in complexity science [bak1996nature, dobson2007complex], we developed a methodology to quantify how system criticality evolves over time:
@@
This approach allowed us to identify critical slowing down and critical fluctuations—hallmark signatures of systems approaching phase transitions [scheffer2009early]. In the context of power grid reliability, a phase transition represents the shift from a regime where outages remain localized and manageable to one where they propagate throughout the system in cascading failures with potentially catastrophic consequences. By tracking the critical exponent’s trajectory relative to the theoretical threshold ($\alpha=1$), we could observe the system’s progression toward criticality, characterized by power-law distributions with increasingly heavier tails.
@@ 2.5 Critical Transition Detection
Complex systems, including power grids, can experience abrupt shifts from one stable state to another, known as critical transitions [scheffer2009early, dakos2015resilience]. The core challenge in complexity science is identifying early warning signals before these transitions occur. Our approach addresses this challenge by analyzing the critical exponent $\alpha$ derived from the power-law distribution of outage sizes.
In SOC theory, the critical exponent $\alpha$ determines the probability distribution of event magnitudes. When $\alpha$ is large, the probability of large-scale events drops rapidly, making them extremely rare. Conversely, when $\alpha$ is small, the decline in probability is more gradual, making larger events relatively more probable.
The value $\alpha=1$ represents a theoretically significant threshold. For a power-law distribution $P(x)\propto x^{-\alpha}$, the normalization condition requires:
$\int_{x_{min}}^{\infty}P(x)\,dx\propto\int_{x_{min}}^{\infty}x^{-\alpha}\,dx=\frac{x^{1-\alpha}}{1-\alpha}\Big|_{x_{min}}^{\infty}$ (3)
This integral converges only when $\alpha>1$. For $\alpha\leq 1$, the integral diverges, meaning the probability mass becomes unbounded toward large events. In practical terms, when $\alpha$ falls below 1, the system enters a regime where extreme events are no longer statistically “rare”—they increasingly dominate the distribution and become inevitable as observation windows extend.
@@
  • Crossings of the $\alpha=1$ threshold, signaling transitions into regimes of heightened vulnerability
  • The duration of periods when the system remains below this threshold
  • The depth of excursion below the threshold (how far $\alpha$ drops below 1)
We hypothesize that these three aspects—threshold crossings, persistence, and excursion depth—together characterize the vulnerability of power systems to catastrophic failures. This framework provides a foundation for interpreting the critical exponent’s behavior in relation to system-wide risk and offers a quantitative approach to identifying periods of heightened vulnerability before catastrophic failures materialize.
The February 2021 Texas power crisis provides a unique case study for testing these predictions. By analyzing how the critical exponent evolved relative to the $\alpha=1$ threshold before, during, and after this well-documented catastrophic failure, we can evaluate whether threshold monitoring provides actionable early warning signals for cascading failures in power systems.